Solution of (D³ - 2D² - 3D)y = 0


A) y = C1.₁e^{x} + C₂е^{-x }+ С3.e^{3x} B) y = C1e^x + C₂е^{-x} + С3.e^{-3x}


C) y = C1+C₂е^{-x} + С3.e^{3x}


D) y = C₁+C₂е^{-x} + С3e^{-3x}


1
Expert's answer
2021-08-18T09:51:01-0400

"\left(D^3-2D^2-3D\right)y=0\\
\mathrm{A\:linear\:homogeneous\:ODE\:with\:constant\:coefficients\:has\:the\:form\:of}\:\left(a_nD^n+...+a_1D+a_0\right)y=0\\
D=0,\:D=-1,\:D=3\\
\mathrm{For\:non\:repeated\:real\:roots\:}D_1,\:D_2,\:...,\:D_n\mathrm{,\:the\:general\:solution\:takes\:the\:form:}\\
y=c_1e^{D_1\:x}+c_2e^{D_2\:x}+...+c_ne^{D_n\:x}\\
c_1e^0+c_2e^{-x}+c_3e^{3x}\\
y=c_1+c_2e^{-x}+c_3e^{3x}"